### v1

# Spectrum Efficiency and Processing Latency Trade-offs in Panel-Based LIS

Abstract
The next generation wireless systems will face stringent new requirements, including ultra-low latency, high data rates and enhanced reliability. Large Intelligent Surfaces, is one proposed solution that has the potential to solve these high demands. The real-life deployment of such systems involves different design considerations with non-trivial trade-offs. This paper investigates the trade-off between spectral efficiency and processing latency, considering different antenna distribution schemes and detection algorithms. A latency model for the physical layer processing has been developed with realistic hardware parameters. The simulation results using an in-door environment show that distributing the antennas throughout the scenario improves the overall reliability, while the impact on the latency is limited when zero-forcing (ZF) detection is used. On the other hand, changing the detection algorithm to maximum-ratio combining (MRC) may reduce the latency significantly, even if a larger number of antennas are needed to achieve a similar spectrum efficiency as ZF detection.
Keywords – 6G, Panel-Based Large Intelligent Surface (LIS), Distributed Massive MIMO, Decentralized Processing, Latency
I Introduction
Distributed multiple-input multiple-output (MIMO) systems are expected to be a part of the up-and-coming wireless systems that can deliver low-latency, high data-rates and reliable communication. One practical implementation of these types of systems is envisioned through the concept of panel-based large intelligent surfaces [1], [2]. One surface is comprised of a large array of antennas, forming a spatially distinct panel, and multiple panels are jointly serving a number of user equipments. Each panel possesses internal processing resources, as well as inter-panel communication capabilities. As the number of antennas increase, so does the amount of data and compute, affecting the total latency of the system, as well as the interconnection bandwidth between the antennas and the central processing unit (CPU). The decentralized approach possible with LIS allows for partially distributing the computations locally, overcoming some of the interconnection bottlenecks [2].
Distributed architectures and algorithms for MIMO [3], [4] and specifically LIS [2], [5], have been suggested and analyzed. In [2], a latency investigation for a panelized LIS system was conducted, with proposed hardware latency models, though without measured latency data from actual hardware. Obtaining real latency numbers from implemented digital signal processing blocks brings valuable insights, especially for understanding practical trade-offs in realistic implementations. These measurements are crucial for defining standards in emerging wireless systems, as they reflect the performance needs of real-world use cases.
For certain linear detection algorithms in MIMO systems, while spectral efficiency improves with the amount of antennas [6], there might still be a large hardware latency overhead, due to the nature and amount of computations that scale with both the number of antennas and the number of users. When comparing linear MIMO receivers, their performance heavily depends on the considered scenario. It is known that zero-forcing (ZF)[7] completely eliminates multi-stream interference at the expense of noise enhancement whereas maximum ratio combining (MRC) does not handle the interference very well, but works better in low signal-to-noise ratio (SNR) scenarios.
This work investigates the trade-offs between performance and latency of the physical layer within a wireless system, when scaling up a panel-based LIS, either by adding more panels or further distributing a fixed amount of antennas. We consider the linear processing schemes presented above and we present a possible algorithmic design when distributing the computing workload locally. Our analysis is based on state-of-the-art implementations of specific hardware for massive MIMO linear equalizers. Moreover, we consider the influence of the radio frame structure on the processing latency, when the system is serving randomly-placed, spatially-multiplexed users.
II System model
Let us consider an orthogonal frequency domain multiplexing (OFDM)-based system where $K$ single antenna UEs transmit in the UL to a panel-based LIS. The panel-based LIS consists of $P$ LIS-panels, each containing $N$ antennas, so that the whole LIS contains a total of $M=NP$ antennas. The vector of symbols received at LIS-panel $p\in\{1,\dots,P\}$ within a given subcarrier may be expressed as
$\boldsymbol{y}_{p}=\boldsymbol{H}_{p}\boldsymbol{s}+\boldsymbol{n}_{p},$ (1)
where $\boldsymbol{H}_{p}$ is the $N\times K$ channel matrix, $\boldsymbol{s}$ is the $K\times 1$ vector of symbols transmitted by the UEs and $\boldsymbol{n}_{p}\sim\mathcal{CN}(0,N_{0}\mathbf{I}_{n})$ corresponds to additive white Gaussian noise (AWGN). Considering the whole LIS, we may express the complete received vector as
$\boldsymbol{y}=\boldsymbol{H}\boldsymbol{s}+\boldsymbol{n},$ (2)
where we defined $\boldsymbol{y}\triangleq[\boldsymbol{y}_{1}^{\mathrm{T}},\dots,\boldsymbol{y}_{P}^{\mathrm{T}}]^{\mathrm{T}}$, $\boldsymbol{H}\triangleq[\boldsymbol{H}_{1}^{\mathrm{T}},\dots,\boldsymbol{H}_{P}^{\mathrm{T}}]^{\mathrm{T}}$, and $\boldsymbol{n}\triangleq[\boldsymbol{n}_{1}^{\mathrm{T}},\dots,\boldsymbol{n}_{P}^{\mathrm{T}}]^{\mathrm{T}}$.
Figure 1: Overview of a general architecture of the processing blocks for panel-based LISs connected in a daisy-chain fashion.
II-A Motivation for decentralized processing
Assuming that all the LIS-panels are connected to a CPU, and that this CPU has access to $\boldsymbol{y}$, as well as to an estimate of the whole $\boldsymbol{H}$, [^1] In order to have the estimate of $\boldsymbol{H}$ at the CPU either channel estimation is performed at this CPU, or the LIS-panels perform channel estimation and forward their estimates of $\boldsymbol{H}_{p}$ $\forall p$ to the CPU. we may coherently equalize the complete received vector by employing common MIMO linear processing schemes. [^2] Linear processing schemes, achieve near-optimal performance as the number of base station (BS) antennas increase [6]. Two useful examples are MRC and ZF [7], whose equalization matrices are given by
$\boldsymbol{W}_{\mathrm{MRC}}=\boldsymbol{H}^{\mathrm{H}},$ (3a)
$\boldsymbol{W}_{\mathrm{ZF}}=(\boldsymbol{H}^{\mathrm{H}}\boldsymbol{H})^{-1}\boldsymbol{H}^{\mathrm{H}},$ (3b)
where the post-processed vector $\boldsymbol{z}=\boldsymbol{W}\boldsymbol{y}$ would be then used to independently decode the data from each user.
An important limitation of centralized processing is that the estimate of the whole matrix $\boldsymbol{H}$, as well as the complete received vector $\boldsymbol{y}$ have to be known at the CPU. This requires sending raw data from the LIS-panels to the CPU, leading to interconnection bandwidths scaling with $M$, and then performing computations at the CPU whose complexity would exhibit the same scaling. A key assumption in systems extending the massive MIMO concept, as LIS, is to have $M\gg K$ [6, 1]. In this case, centralized processing becomes unscalable from the excessive interconnection bandwidth and processing complexity requirements [8]. An alternative is to alleviate those requirements by performing part of the processing parallelly at the LIS-panels using decentralized processing approaches, as will be next explored.
We consider decentralized processing approaches where the resulting interconnection bandwidth to a CPU, as well as the respective processing complexity, scale with the number of UEs $K$ instead of with the total number of antennas $M$ [^3] As argued in [8], the data and resources associated to $K$ simultaneous UEs inevitably scale with $K$ unless part of the end data is to be ignored. To this end, each LIS-panel $p$ is associated to a local processing unit (LPU) which can estimate the local channel $\boldsymbol{H}_{p}$ and perform simple computations based on this estimate. For simplicity we assume that perfect estimates of $\boldsymbol{H}_{p}$ are obtained, since estimation errors would essentially compromise performance in the same way as for a centralized processing scheme. Moreover, we consider that the LIS-panels are sequentially connected via unidirectional links, i.e., forming a Daisy chain topology as considered in [9]. Each link may then be used to share data between panels, where each link would at most support interconnection bandwidth scaling with $K$. The considered architecture is illustrated in Fig. 1.
III Implementation Considerations
III-A Algorithmic design
If we examine the linear processing schemes given in (3), which assume centralized processing, we may note that the resulting post-processed vectors may be expressed in terms of the local channels by
$\boldsymbol{z}_{\mathrm{MRC}}$ $=\sum_{p=1}^{P}\boldsymbol{H}_{p}^{\mathrm{H}}\boldsymbol{y}_{p}$ (4a)
$=\sum_{p=1}^{P}\boldsymbol{z}_{\mathrm{MRC},p},$
$\boldsymbol{z}_{\mathrm{ZF}}$ $=\boldsymbol{G}^{-1}\boldsymbol{z}_{\mathrm{MRC}}$ (4b)
$=\bigg{(}\sum_{p=1}^{P}\boldsymbol{G}_{p}\bigg{)}^{-1}\boldsymbol{z}_{\mathrm{MRC}},$
where $\boldsymbol{z}_{\mathrm{MRC},p}$ is the local MRC-post-processed vector and $\boldsymbol{G}_{p}=\boldsymbol{H}^{\mathrm{H}}_{p}\boldsymbol{H}_{p}$ is the local Gramian at LIS panel $p$. Hence, we can achieve the same performance as with centralized processing by simply computing at each panel the local MRC-post-processed vector and the local Gramian, and aggregating these throughout the network. Note that $\boldsymbol{z}_{\mathrm{MRC},p}$ is a $K\times 1$ vector and $\boldsymbol{G}_{p}$ is an $K\times K$, so their dimensions scale with the number of UEs, but not with the total number of antennas.
Figure 2: A diagram showing the latency of the system and its different sources.
Considering the Daisy chain interconnection topology, the aggregation of $\boldsymbol{z}_{\mathrm{MRC},p}$ and $\boldsymbol{G}_{p}$ can be done sequentially. Assuming that the LIS-panels are ordered according to their position in the Daisy chain, the starting LIS-panel $1$ would then share directly $\boldsymbol{z}_{\mathrm{MRC},1}$ and $\boldsymbol{G}_{1}$ to LIS-panel $2$, which would add them with its respective contributions $\boldsymbol{z}_{\mathrm{MRC},2}$ and $\boldsymbol{G}_{2}$ and share the results to the next LIS-panel. This would go on sequentially until the final LIS-panel $P$ is reached. After adding its contributions, LIS-panel $P$ would thus have a complete estimate of both $\boldsymbol{z}_{\mathrm{MRC}}$ and $\boldsymbol{G}$, which may be used to perform ZF equalization (by inverting $\boldsymbol{G}$ and multiplying it to $\boldsymbol{z}_{\mathrm{MRC}}$), while the data shared through each connection would scale quadratically with $K$. On the other hand, if we would restrict the processing to MRC, there would be no need for sharing and aggregating the $\boldsymbol{G}_{p}$ matrices, so the data shared through each connection would only scale linearly with $K$. However, we should note that the $\boldsymbol{z}_{\mathrm{MRC},p}$ vectors are directly related to transmitted symbols, while the $\boldsymbol{G}_{p}$ matrices are only related to the channel used for processing. Thus, for large coherence bandwidths we could reduce the scaling of the data required to share the aggregated $\boldsymbol{G}_{p}$ matrices through each connection. For example, if the coherence bandwidth is large enough so that we can perform channel estimation with a single time-slot, i.e., by having the UEs transmit their pilots over non-overlapping subcarriers, both MRC and ZF would require interconnection bandwidths scaling linearly with the number of UEs. In any case, the total number of sequential unidirectional transmissions needed to implement the previous equalizers in a Daisy chain decentralized architecture would be $P-1$.
III-B Reference processing architecture
We consider the hardware architecture in Fig. 1 for $P$ panels, each having $N$ antennas. Each antenna corresponds to a processing chain within the LPUs. The chain begins with the Radio Frequency (RF) stage, which amplifies, filters the analog signal and converts it to the digital domain. It then moves to the digital front-end, where the signal is downsampled, filtered, and transformed via a fast fourier transform (FFT) to obtain the frequency-domain signal. The processed signals from all antennas within one panel are aggregated into a matrix form in order to obtain the local MRC-post-processed vector and the local Gramian, given in (4a) and (4b), respectively. The data is then prepared to be transferred between panels through the fronthaul infrastructure. For panels $p>1$, the local contributions are summed with the incoming data from the preceding panels in the Daisy chain. The final panel in the chain acts as both a LPU for its antennas, and as a CPU for the fully aggregated data. For MRC, the equalization is done after the aggregation, and for ZF matrix inversion for the Gramian and subsequent multiplication with the MRC vector is performed, (4b), in the CPU.
The local and central MIMO processing can be performed with different types of hardware architectures, ranging from application-specific integrated circuits to more generalized architectures such as a general purpose processors. It is important to note that performance metrics could vary greatly depending on the chosen architecture. For this paper, we consider one of the processing units implemented in [10]: a vector core, part of an application-specific instruction set processor (asip). It can be leveraged for both the local and central MIMO processing, and it is specifically tailored for MIMO equalization operations. An asip is higher performing than a standard CPU, but more flexible than an ASIC, and it can be programmed to execute the processing for different amount of UEs and antennas[10].
TABLE I: Latency measured for the asip[10] considered with 16 antennas per panel and 16 or 128 UEs.
Operation Latency (clock cycles) Complexity
K=16 K=128 order
$\boldsymbol{H}=\boldsymbol{y}\boldsymbol{S}^{-1}_{\text{Pilot}}$ 83 531 $\mathcal{O}(KN)$
$\boldsymbol{G}=\boldsymbol{H}^{\mathrm{H}}\boldsymbol{H}$ 946 58k $\mathcal{O}(K^{2}N)$
$\boldsymbol{G}^{-1}$ 1817 890k $\mathcal{O}(K^{3})$
$\boldsymbol{z}_{\mathrm{MRC}}=\boldsymbol{H}^{\mathrm{H}}\boldsymbol{y}$ 81 460 $\mathcal{O}(KN)$
$\boldsymbol{z}_{\mathrm{ZF}}=\boldsymbol{G}^{-1}\boldsymbol{z}_{\mathrm{MRC}}$ 81 3300 $\mathcal{O}(K^{2})$
The vector core consists of 16 single instruction, multiple data (SIMD) lanes and supports different matrix and vector sizes. The considered processing times for different matrix and vector operations are shown in Table I with $N=16$ and $K=16$ or $K=128$, together with the complexities order of each operation. The inverse of the Gramian is calculated using extended QRD, due to its high performance on this specific vector core.
The remaining digital logic blocks are implemented on the AMD ZCU216 FPGA to extract the latency numbers from real hardware. The fronthaul is Ethernet based, and routed directly between panels. The data aggregation is integrated with the Ethernet so that the aggregation can begin immediately when the first data is received. This removes the need for long buffers in relation to the Ethernet. For all hardware units, a clock frequency of $150$MHz is assumed.
IV Latency Analysis
IV-A Latency Model
In Fig. 2 a break-down of the system latency is displayed. We define the latency as the time starting when the UE has new data to transmit and ending when the data for the first subcarrier has been detected by the BS, using either ZF or MRC algorithms. The latency can be expressed as
$\tau=\tau_{\mathrm{wait}}+\tau_{\mathrm{LPU}}+\tau_{\mathrm{fronthaul}}+\tau_{\mathrm{CPU}}.$ (5)
It consists of two main sources, the latency due to the frame structure $\tau_{\mathrm{wait}}$ and the latency due to the physical layer processing in the BS. Based on the architecture in Fig. 1, the BS processing latency has three parts $\tau_{\mathrm{LPU}}$, $\tau_{\mathrm{fronthaul}}$, and $\tau_{\mathrm{CPU}}$, which are the latency in the BS cause by the LPU, front-haul and CPU respectively.
A few assumptions are made regarding the system throughput to simplify latency calculations. It is assumed that there are enough hardware resources to match the data throughput of both in-panel processing and inter-panel data transfer. These assumptions gives the liberty to measure the individual latency of every source in the system without taking the throughput into account (e.g., the extra latency due to buffering can be neglected). There are also a few latency sources that we disregard in the analysis, as they are too small to have any impact. The latency due to the RF chain in the LPU is known to be very small[11] and, as it is done in parallel for all antennas, it can safely be ignored. Also, as an indoor scenario is considered, the signal propagation, both the one over air between the UE and BS, and the one between the panels, are not taken into account.
When computing $\tau_{\mathrm{wait}}$, we calculate the longest time from when the UE has new data to transmit until there is an available data (ULD) slot. This latency will depend on both the frame structure and the OFDM transmission time. In this analysis, we consider a fixed frame structure as presented in Fig. 2. It is composed of a total of 7 OFDM slots, consisting of UL and downlink (DL) transmission, as well as guardband. The OFDM transmission time, including the cyclic prefix, is $19\mu s$, and a subcarrier spacing of 60kHz is assumed. The highest $\tau_{\mathrm{wait}}$ happen when the UE gets new data to transmit as the second ULD symbol has just started transmitting, as it then has to wait for six OFDM symbols before the next available ULD slot. This makes $\tau_{\mathrm{wait}}=133\mu s$. Due to this long wait at the UE, any processing necessary at the UE side can safely be disregarded, as it will be able to finish within this time frame.
Figure 3: The latency results, in both percentage and actual values, depending on the amount of UEs, for a system with 16 antennas per panel and 127 panels, for both ZF and MRC.
For LPU the latency, $\tau_{\mathrm{LPU}}$, is divided into two sources. The latency from the implemented digital front end and FFT block has a latency of $25\mu s$. It is quite high due to the fact that the full OFDM symbol must be received before the computations can be completed. The latency from the local MIMO processing is not as straightforward to determine, due to the fact that this latency will depend on both $N$ and $K$. Also, both the processing for the UL data symbol and the UL pilot need to be taken into account, as the calculations for the UL data cannot start until the channel estimate for the relevant subcarrier is available. In ZF, the latency for the local Gramian computation also needs to be taken into account, as it needs to be transmitted through the fronthaul architecture before the local MRC vector. The complexity of the relevant calculations are shown in Table I.
The latency of the Ethernet-based fronthaul, $\tau_{\mathrm{fronthaul}}$, grows linearly with the amount of panels according to
$\tau_{\mathrm{fronthaul}}=0.87\times(P-1),$ (6)
where $0.87$ is the cost for every interconnection step expressed in $\mu s$.
For MRC $\tau_{\mathrm{CPU}}$ will be zero, but for ZF both the latency of the Gramian inversion and matrix multiplication needs to be taken into account. The complexity of these calculations is also included in Table I.
IV-B Latency Results
Figure 4: Example scenarios used in the spectral efficiency simulations. $P=1$ with a $32\times 32$ panel (left) and $P=32$ with $8\times 4$ panels (right).
In Fig. 3 the latency for a system with $N=16$ and $P=128$ is shown for both ZF and MRC. When the number of UEs is relatively small, the latency distribution of different components are similar, with almost half being due to the frame structure, roughly a third due to the Ethernet-based fronthaul and the rest caused by the processing. However, when the amount of UEs grows, the two approaches lead to completely different latency ratios. For MRC, increasing the amount of UEss does not affect the latency much, as the needed computations only grow linearly with $K$ and are quite simple to execute. For ZF though, the processing latency is completely dominant for a large amount of UEs as the local processing grows quadratically with $K$ and the central processing grows cubically with $K$.
To decrease the latency of a distributed massive multiple-input multiple-output (DMIMO) system using ZF and which is designed to support a large amount of UEs, further optimizations clearly need to be done. A simple solution would be to further optimize the processing, though that would likely lead to higher resource and energy costs. Another interesting avenue that could be investigated is the potential of panel segmentation, where only a subset of panels would serve certain UEs. This would lead to smaller matrix sizes and therefore simplifying the processing, but it would cause new unique problems.
V Latency Versus Spectral Efficiency
One of the main performance metrics in any wireless communication scenario is the data rate supported by the system. In the considered setting, we are specially interested in the data rates experienced by the different UEs when transmitting their UL data, and how they interplay with the arrangement of the LIS into panels, as well as with the resulting physical layer latency. To this end, we consider an industrial indoor scenario, where the channel matrix follows the Rician fading model [7]
$\boldsymbol{H}=\sqrt{\frac{K_{\mathrm{F}}}{1+K_{\mathrm{F}}}}\boldsymbol{H}_{\mathrm{LoS}}+\frac{1}{\sqrt{1+K_{\mathrm{F}}}}\boldsymbol{H}_{\mathrm{R}},$ (7)
where $K_{\mathrm{F}}$ is the Rician factor, $\boldsymbol{H}_{\mathrm{LoS}}$ is the line of sight (LoS) channel between the UEs and the whole LIS, and $\boldsymbol{H}_{\mathrm{R}}$ follows the IID Rayleigh fading model where the average power for each entry equals the power of the respective LoS contribution in $\boldsymbol{H}_{\mathrm{LoS}}$. The parameters of the scenario are summarized in Table II. Note that, since we consider LIS scenarios with a large number of panels deployed throughout compact indoor regions, we should take into account near-field effects when computing the LoS contributions [12]. Specifically, we consider that only the perpendicular incident power is absorbed by each antenna element, and we construct the array steering vectors using the spherical propagation model. Moreover, we normalize the channel power such that the SNR value from Table II corresponds to the average receive SNR considering the whole LIS, where the average is taken over the whole set of simulated scenarios being compared.
TABLE II: Scenario parameters.
Parameter Value
Frequency $3.2$ GHz
Antenna spacing $\lambda/2$
Rician factor $5$ dB
Parameter Value
SNR $10$ dB
Room length $30$ m
Room width $20$ m
We begin by analyzing the effect of distributing a fixed number of antennas into a varying number of LIS-panels. Specifically, we consider deploying $M=1024$ antennas into equidistant panels along the two walls extending through the length of the room. Note that, in all cases (except when all the antennas are co-located), each panel is facing another panel. The number of panels is incremented from $P=1$ to $P=256$ by powers of 2, while for each $P$ value the respective panels are either squared or have twice as many elements along their width than along their height. There are $K=128$ UEs transmitting equal power, and these are randomly located throughout a plane at a minimum distance of $10\lambda$ from the walls. The LIS-panels extend from the UEs plane towards the ceiling of the room. An example scenario is given in Fig. 4. All the spectral efficiency results are averaged over $10^{2}$ realizations of the uniform UE placements, each including $10^{3}$ realizations of the Rayleigh fading channel component. Fig. 5 plots the spectral efficiencies with respect to the number of panels for the described scenarios. We can see that distributing the antennas throughout the scenario into smaller LIS-panels favors user fairness, reducing potential outages for both ZF and MRC processing. We believe that this is a strong motivation for favoring more distributed antenna deployments since it increases the overall reliability. On the other hand, as seen in the previous section, in dense user deployments the latency values for the ZF case are dominated by the central processing required to perform matrix inversion, so the most distributed implementation has marginal latency increase over having all the antennas co-located in one panel. In the MRC case, the fronthaul latency associated to sharing the data sequentially between panels still has noticeable impact, leading to a latency-performance trade-off which should be further analyzed under different settings. Nevertheless, there is a major latency reduction with respect to ZF, at the cost of some degradation in average user spectral efficiency.
Figure 5: User spectral efficiencies versus number of panels in logarithmic scale for fixed total number of antennas.
Figure 6: Average user spectral efficiency versus number of $4\times 4$ panels in logarithmic scale. For each point, the computed latency value is included.
We now consider the scaling of the spectral efficiencies when deploying varying number $P$ of equally sized square panels with $N=16$ antennas. Fig. 6 shows the average user spectral efficiency with respect to the number of panels for the indoor scenario previously described. Aside from the fixed panel size, the same scenario considerations apply as for the results in Fig. 6. As one may expect, the average user spectral efficiency increases logarithmically with the number of panels for high enough $P$. This may be justified by the extra array gain associated to having more panels receiving the transmitted power. We have also included for comparison the latency values for each panel deployment. Interestingly, the spectral efficiency achieved by MRC is comparable to that achieved by ZF with half the LIS-panels, which still requires a latency close to $30$ times higher. Thus, we can conclude that, in the considered scenario, MRC is much more effective at exploiting the trade-off between latency and spectral efficiency.
VI Conclusion
We have studied the latency in a panel-based LIS indoor scenario with different panel deployments. We presented algorithms for performing decentralized MRC and ZF processing without any loss with respect to the centralized counterpart. We also computed latency values based on real implementations for the different schemes, which show that, as the number of users grows, the ZF latency is quickly dominated by the central processing required to invert the channels. Finally, we compared the spectral efficiency achieved with different deployments, where we observed that having antennas more distributed increases the system reliability, and comes at essentially no latency cost when using ZF. However, MRC seems to attain a better trade-off between latency and performance.

### v3

@@
 This paper investigates the trade-off between spectral efficiency and processing latency, considering different antenna distribution schemes and detection algorithms.
-A latency model for the physical layer processing has been developed with realistic hardware parameters.
-The simulation results using an in-door environment show that distributing the antennas throughout the scenario improves the overall reliability, while the impact on the latency is limited when zero-forcing (ZF) detection is used.
-On the other hand, changing the detection algorithm to maximum-ratio combining (MRC) may reduce the latency significantly, even if a larger number of antennas are needed to achieve a similar spectrum efficiency as ZF detection.
+A latency model for the physical layer processing has been developed, using real FPGA and application-specific instruction processor (ASIP) hardware implementation results.
+Simulation results using an indoor environment show that distributing antennas throughout the scenario improves overall reliability, while the impact from this on latency is limited both when using zero-forcing (ZF) and Minimum Mean Square Error (MMSE) detection.
+Changing the detection algorithm to maximum-ratio combining (MRC) from ZF or MMSE, however, reduces the latency significantly, even if a larger number of antennas are needed to achieve a similar spectrum efficiency.
 Keywords – 6G, Panel-Based Large Intelligent Surface (LIS), Distributed Massive MIMO, Decentralized Processing, Latency
@@
 Distributed multiple-input multiple-output (MIMO) systems are expected to be a part of the up-and-coming wireless systems that can deliver low-latency, high data-rates and reliable communication.
-One practical implementation of these types of systems is envisioned through the concept of panel-based large intelligent surfaces [1], [2].
-One surface is comprised of a large array of antennas, forming a spatially distinct panel, and multiple panels are jointly serving a number of user equipments.
+One implementation of such a systems is envisioned through the concept of panel-based active large intelligent surface (LIS) with fully digital beamforming, [1], [2].
+In this case the LIS consist of a large number of antenna elements with individual coherent transceiver chains which, in this case, are composed of a number of connected panels, each constituting a sub-array.
+In this configuration, the panels are jointly serving a number of user equipments.
 Each panel possesses internal processing resources, as well as inter-panel communication capabilities.
-As the number of antennas increase, so does the amount of data and compute, affecting the total latency of the system, as well as the interconnection bandwidth between the antennas and the central processing unit (CPU).
-The decentralized approach possible with LIS allows for partially distributing the computations locally, overcoming some of the interconnection bottlenecks [2].
+As the number of antennas increase, so does the amount of data and compute, affecting the total latency of the system and the interconnection bandwidth between antennas and the central processing unit (CPU).
+Decentralized processing approaches available for LIS allow overcoming some of the interconnection bottlenecks by distributing the computations throughout local processing units [2], [3].
 Distributed architectures and algorithms for MIMO [4], [5] and specifically LIS [2], [3], have been suggested and analyzed.
@@
 Obtaining real latency numbers from implemented digital signal processing blocks brings valuable insights, especially for understanding practical trade-offs in realistic implementations.
-These measurements are crucial for defining standards in emerging wireless systems, as they reflect the performance needs of real-world use cases.
-For certain linear detection algorithms in MIMO systems, while spectral efficiency improves with the amount of antennas [6], there might still be a large hardware latency overhead, due to the nature and amount of computations that scale with both the number of antennas and the number of users.
-When comparing linear MIMO receivers, their performance heavily depends on the considered scenario.
-It is known that zero-forcing (ZF)[7] completely eliminates multi-stream interference at the expense of noise enhancement whereas maximum ratio combining (MRC) does not handle the interference very well, but works better in low signal-to-noise ratio (SNR) scenarios.
-This work investigates the trade-offs between performance and latency of the physical layer within a wireless system, when scaling up a panel-based LIS, either by adding more panels or further distributing a fixed amount of antennas.
-We consider the linear processing schemes presented above and we present a possible algorithmic design when distributing the computing workload locally.
+These measurements are crucial for defining standards in emerging wireless systems, as they have to meet the performance needs of real-world use cases.
+[…]
+This work investigates the trade-offs between performance and latency of the physical layer within a panel-based LIS scenario where we scale the LIS by adding more panels or further distributing a fixed amount of antennas.
+We consider the linear processing schemes presented above and we present a possible algorithmic design when distributing the total computing workload onto local panels.
 Our analysis is based on state-of-the-art implementations of specific hardware for massive MIMO linear equalizers.
-Moreover, we consider the influence of the radio frame structure on the processing latency, when the system is serving randomly-placed, spatially-multiplexed users.
+Moreover, we consider the influence of the radio frame structure on the processing latency when the system is serving randomly-placed spatially-multiplexed users.
+Figure 1: A diagram showing the latency of the system and its different sources.
 II System model
-Let us consider an orthogonal frequency domain multiplexing (OFDM)-based system where $K$ single antenna UEs transmit in the UL to a panel-based LIS.
+Let us consider an orthogonal frequency-division multiplexing (OFDM)-based system where $K$ single antenna UEs transmit in the UL to a panel-based LIS.
 The panel-based LIS consists of $P$ LIS-panels, each containing $N$ antennas, so that the whole LIS contains a total of $M=NP$ antennas.
@@
 $\boldsymbol{y}_{p}=\boldsymbol{H}_{p}\boldsymbol{s}+\boldsymbol{n}_{p},$ (1)
-where $\boldsymbol{H}_{p}$ is the $N\times K$ channel matrix, $\boldsymbol{s}$ is the $K\times 1$ vector of symbols transmitted by the UEs and $\boldsymbol{n}_{p}\sim\mathcal{CN}(0,N_{0}\mathbf{I}_{n})$ corresponds to additive white Gaussian noise (AWGN).
-Considering the whole LIS, we may express the complete received vector as
+where $\boldsymbol{H}_{p}$ is the $N\times K$ channel matrix, $\boldsymbol{s}$ is the $K\times 1$ vector of symbols transmitted by the UEs and $\boldsymbol{n}_{p}\sim\mathcal{CN}(0,N_{0}\mathbf{I}_{n})$ corresponds to additive white Gaussian noise.
+Considering the whole LIS, we may express the complete received vector across all P panels as
 $\boldsymbol{y}=\boldsymbol{H}\boldsymbol{s}+\boldsymbol{n},$ (2)
@@
 II-A Motivation for decentralized processing
-Assuming that all the LIS-panels are connected to a CPU, and that this CPU has access to $\boldsymbol{y}$, as well as to an estimate of the whole $\boldsymbol{H}$, [^1] In order to have the estimate of $\boldsymbol{H}$ at the CPU either channel estimation is performed at this CPU, or the LIS-panels perform channel estimation and forward their estimates of $\boldsymbol{H}_{p}$ $\forall p$ to the CPU.
+Assuming that all LIS-panels are connected to a CPU, and that this CPU has access to $\boldsymbol{y}$, as well as to an estimate of the whole $\boldsymbol{H}$, [^1] In order to have the estimate of $\boldsymbol{H}$ at the CPU either channel estimation is performed at this CPU, or the LIS-panels perform channel estimation and forward their estimates of $\boldsymbol{H}_{p}$ $\forall p$ to the CPU.
 we may coherently equalize the complete received vector by employing common MIMO linear processing schemes.
-[^2] Linear processing schemes, achieve near-optimal performance as the number of base station (BS) antennas increase [6].
-Two useful examples are MRC and ZF [7], whose equalization matrices are given by
+[^2] Linear processing schemes achieve near-optimal performance as the number of base station antennas increase [6].
+Three common linear equalizers are MRC, ZF and minimum mean squared error (MMSE) [7], whose equalization matrices are given by
 $\boldsymbol{W}_{\mathrm{MRC}}=\boldsymbol{H}^{\mathrm{H}},$ (3a)
-$\boldsymbol{W}_{\mathrm{ZF}}=(\boldsymbol{H}^{\mathrm{H}}\boldsymbol{H})^{-1}\boldsymbol{H}^{\mathrm{H}},$ (3b)
-where the post-processed vector $\boldsymbol{z}=\boldsymbol{W}\boldsymbol{y}$ would be then used to independently decode the data from each user.
-An important limitation of centralized processing is that the estimate of the whole matrix $\boldsymbol{H}$, as well as the complete received vector $\boldsymbol{y}$ have to be known at the CPU.
-This requires sending raw data from the LIS-panels to the CPU, leading to interconnection bandwidths scaling with $M$, and then performing computations at the CPU whose complexity would exhibit the same scaling.
-A key assumption in systems extending the massive MIMO concept, as LIS, is to have $M\gg K$ [6, 1].
-In this case, centralized processing becomes unscalable from the excessive interconnection bandwidth and processing complexity requirements [8].
-An alternative is to alleviate those requirements by performing part of the processing parallelly at the LIS-panels using decentralized processing approaches, as will be next explored.
-We consider decentralized processing approaches where the resulting interconnection bandwidth to a CPU, as well as the respective processing complexity, scale with the number of UEs $K$ instead of with the total number of antennas $M$ [^3] As argued in [8], the data and resources associated to $K$ simultaneous UEs inevitably scale with $K$ unless part of the end data is to be ignored.
-To this end, each LIS-panel $p$ is associated to a local processing unit (LPU) which can estimate the local channel $\boldsymbol{H}_{p}$ and perform simple computations based on this estimate.
+$\boldsymbol{W}_{\mathrm{ZF}}=(\boldsymbol{H}^{\mathrm{H}}\boldsymbol{H})^{-1}\boldsymbol{H}^{\mathrm{H}},\text{ and}$ (3b)
+$\boldsymbol{W}_{\mathrm{MMSE}}=(\boldsymbol{H}^{\mathrm{H}}\boldsymbol{H}+N_{0}\boldsymbol{I}_{K})^{-1}\boldsymbol{H}^{\mathrm{H}},$ (3c)
+respectively, and the post-processed vector $\boldsymbol{z}=\boldsymbol{W}\boldsymbol{y}$ would be then used to decode user data.
+An important limitation of centralized processing is that both the estimate of the whole matrix $\boldsymbol{H}$ and the complete received vector $\boldsymbol{y}$ have to be known at the CPU.
+This requires sending raw data from the LIS-panels to the CPU, leading to interconnection bandwidths scaling with number of antennas, $M$, and then performing computations at the CPU whose complexity would exhibit the same scaling.
+A key assumption in systems extending the massive MIMO concept, as LIS, is to have $M\gg K$ [6], [1].
+In this case, centralized processing in a distributed system becomes unscalable from the excessive interconnection bandwidth and processing complexity requirements [4].
+An alternative is to alleviate those requirements by performing part of the processing in parallel at the LIS-panels using decentralized processing approaches, as explored next.
+We consider decentralized processing approaches where the resulting interconnection bandwidth to a CPU, as well as the respective processing complexity, scale with the number of UEs $K$ instead of with the total number of antennas MM.
+[^3] As argued in [4], the data and resources associated to $K$ simultaneous UEs inevitably scale with $K$ unless part of the end data is to be ignored.
+To this end, each LIS-panel $p$ is associated to a LPU which can estimate the local channel $\boldsymbol{H}_{p}$ and perform simple computations based on this estimate.
 For simplicity we assume that perfect estimates of $\boldsymbol{H}_{p}$ are obtained, since estimation errors would essentially compromise performance in the same way as for a centralized processing scheme.
-Moreover, we consider that the LIS-panels are sequentially connected via unidirectional links, i.e., forming a Daisy chain topology as considered in [9].
-Each link may then be used to share data between panels, where each link would at most support interconnection bandwidth scaling with $K$.
+Moreover, we consider that the LIS-panels are sequentially connected via unidirectional links, forming a Daisy chain topology [8].
+Each link may then be used to share data between panels and would at most support interconnection bandwidth scaling with $K$.
 The considered architecture is illustrated in Fig.
-1.
+2.
 III Implementation Considerations
@@
 If we examine the linear processing schemes given in (3), which assume centralized processing, we may note that the resulting post-processed vectors may be expressed in terms of the local channels by
-$\boldsymbol{z}_{\mathrm{MRC}}$ $=\sum_{p=1}^{P}\boldsymbol{H}_{p}^{\mathrm{H}}\boldsymbol{y}_{p}$ (4a)
-$=\sum_{p=1}^{P}\boldsymbol{z}_{\mathrm{MRC},p},$
-$\boldsymbol{z}_{\mathrm{ZF}}$ $=\boldsymbol{G}^{-1}\boldsymbol{z}_{\mathrm{MRC}}$ (4b)
-$=\bigg{(}\sum_{p=1}^{P}\boldsymbol{G}_{p}\bigg{)}^{-1}\boldsymbol{z}_{\mathrm{MRC}},$
+$\boldsymbol{z}_{\mathrm{MRC}}$ $=\sum_{p=1}^{P}\boldsymbol{H}_{p}^{\mathrm{H}}\boldsymbol{y}_{p}=\sum_{p=1}^{P}\boldsymbol{z}_{\mathrm{MRC},p},$ (4a)
+$\boldsymbol{z}_{\mathrm{ZF}}$ $=\boldsymbol{G}^{-1}\boldsymbol{z}_{\mathrm{MRC}}=\bigg(\sum_{p=1}^{P}\boldsymbol{G}_{p}\bigg)^{-1}\boldsymbol{z}_{\mathrm{MRC}},$ (4b)
+$\boldsymbol{z}_{\mathrm{MMSE}}=\bigg(\sum_{p=1}^{P}\boldsymbol{G}_{p}+N_{0}\boldsymbol{I}_{K}\bigg)^{-1}\boldsymbol{z}_{\mathrm{MRC}},$ (4c)
 where $\boldsymbol{z}_{\mathrm{MRC},p}$ is the local MRC-post-processed vector and $\boldsymbol{G}_{p}=\boldsymbol{H}^{\mathrm{H}}_{p}\boldsymbol{H}_{p}$ is the local Gramian at LIS panel $p$.
 Hence, we can achieve the same performance as with centralized processing by simply computing at each panel the local MRC-post-processed vector and the local Gramian, and aggregating these throughout the network.
-Note that $\boldsymbol{z}_{\mathrm{MRC},p}$ is a $K\times 1$ vector and $\boldsymbol{G}_{p}$ is an $K\times K$, so their dimensions scale with the number of UEs, but not with the total number of antennas.
-Figure 2: A diagram showing the latency of the system and its different sources.
+Note that $\boldsymbol{z}_{\mathrm{MRC},p}$ is a $K\times 1$ vector and $\boldsymbol{G}_{p}$ is an $K\times K$ matrix, so their dimensions scale with the number of UEs, but not with the total number of antennas.
 Considering the Daisy chain interconnection topology, the aggregation of $\boldsymbol{z}_{\mathrm{MRC},p}$ and $\boldsymbol{G}_{p}$ can be done sequentially.
-Assuming that the LIS-panels are ordered according to their position in the Daisy chain, the starting LIS-panel $1$ would then share directly $\boldsymbol{z}_{\mathrm{MRC},1}$ and $\boldsymbol{G}_{1}$ to LIS-panel $2$, which would add them with its respective contributions $\boldsymbol{z}_{\mathrm{MRC},2}$ and $\boldsymbol{G}_{2}$ and share the results to the next LIS-panel.
+Assuming that the LIS-panels are numbered according to their position in the Daisy chain, the starting LIS-panel $1$ would then share directly $\boldsymbol{z}_{\mathrm{MRC},1}$ and $\boldsymbol{G}_{1}$ to LIS-panel $2$, which would add them with its respective contributions $\boldsymbol{z}_{\mathrm{MRC},2}$ and $\boldsymbol{G}_{2}$ and share the results to the next LIS-panel.
 This would go on sequentially until the final LIS-panel $P$ is reached.
-After adding its contributions, LIS-panel $P$ would thus have a complete estimate of both $\boldsymbol{z}_{\mathrm{MRC}}$ and $\boldsymbol{G}$, which may be used to perform ZF equalization (by inverting $\boldsymbol{G}$ and multiplying it to $\boldsymbol{z}_{\mathrm{MRC}}$), while the data shared through each connection would scale quadratically with $K$.
-On the other hand, if we would restrict the processing to MRC, there would be no need for sharing and aggregating the $\boldsymbol{G}_{p}$ matrices, so the data shared through each connection would only scale linearly with $K$.
-However, we should note that the $\boldsymbol{z}_{\mathrm{MRC},p}$ vectors are directly related to transmitted symbols, while the $\boldsymbol{G}_{p}$ matrices are only related to the channel used for processing.
-Thus, for large coherence bandwidths we could reduce the scaling of the data required to share the aggregated $\boldsymbol{G}_{p}$ matrices through each connection.
-For example, if the coherence bandwidth is large enough so that we can perform channel estimation with a single time-slot, i.e., by having the UEs transmit their pilots over non-overlapping subcarriers, both MRC and ZF would require interconnection bandwidths scaling linearly with the number of UEs.
-In any case, the total number of sequential unidirectional transmissions needed to implement the previous equalizers in a Daisy chain decentralized architecture would be $P-1$.
+After adding its contributions, LIS-panel $P$ has a complete estimate of both $\boldsymbol{z}_{\mathrm{MRC}}$ and $\boldsymbol{G}$, which may be used to perform ZF or MMSE equalization, i.e., by inverting $\boldsymbol{G}$ (or its MMSE regularized version) and multiplying it to $\boldsymbol{z}_{\mathrm{MRC}}$.
+In this case data shared through each connection would scale quadratically with $K$.
+On the other hand, if we restrict processing to MRC, there is no need for sharing and aggregating the $\boldsymbol{G}_{p}$ matrices, so the data shared through each connection only scales linearly with $K$.
+However, we should note that the $\boldsymbol{z}_{\mathrm{MRC},p}$ vectors are directly related to transmitted symbols, while the $\boldsymbol{G}_{p}$ matrices are only related to the channel.
+Thus, for large coherence bandwidths we can reduce the scaling of the data required to share the aggregated $\boldsymbol{G}_{p}$ matrices through each connection.
+For example, if the coherence bandwidth is large enough so that we can perform channel estimation with a single time-slot, i.e., by having the UEs transmit their pilots over non-overlapping subcarriers, MRC, ZF, and MMSE would all require interconnection bandwidths scaling linearly with the number of UEs.
+In any case, the total number of sequential unidirectional transmissions needed to implement the previous equalizers in a Daisy chain decentralized architecture becomes P1P-1.
 III-B Reference processing architecture
@@
 2 for $P$ panels, each having $N$ antennas.
-Each antenna corresponds to a processing chain within the LPUs.
+Each antenna is connected to a processing chain within the LPUs.
 The chain begins with the Radio Frequency (RF) stage, which amplifies, filters the analog signal and converts it to the digital domain.
 It then moves to the digital front-end, where the signal is downsampled, filtered, and transformed via a fast fourier transform (FFT) to obtain the frequency-domain signal.
-The processed signals from all antennas within one panel are aggregated into a matrix form in order to obtain the local MRC-post-processed vector and the local Gramian, given in (4a) and (4b), respectively.
+The processed signals from all antennas within one panel are aggregated into a matrix form in order to obtain the local MRC-post-processed vector and the local Gramian, given in (4).
 The data is then prepared to be transferred between panels through the fronthaul infrastructure.
 For panels $p>1$, the local contributions are summed with the incoming data from the preceding panels in the Daisy chain.
-The final panel in the chain acts as both a LPU for its antennas, and as a CPU for the fully aggregated data.
-For MRC, the equalization is done after the aggregation, and for ZF matrix inversion for the Gramian and subsequent multiplication with the MRC vector is performed, (4b), in the CPU.
+The final panel in the chain acts both as a LPU for its antennas, and as a CPU for the fully aggregated data.
+For MRC, the equalization is done after the aggregation, and for ZF and MMSE, matrix inversion of the Gramian (or the SNR regularized version for MMSE) and subsequent multiplication with the MRC vector is performed in the CPU.
 The local and central MIMO processing can be performed with different types of hardware architectures, ranging from application-specific integrated circuits to more generalized architectures such as a general purpose processors.
-It is important to note that performance metrics could vary greatly depending on the chosen architecture.
+It is important to note that performance metrics can vary greatly depending on the chosen architecture.
 For this paper, we consider one of the processing units implemented in [9]: a vector core, part of an application-specific instruction set processor (ASIP).
@@
 $\boldsymbol{z}_{\mathrm{ZF}}=\boldsymbol{G}^{-1}\boldsymbol{z}_{\mathrm{MRC}}$ 81 3300 $\mathcal{O}(K^{2})$
-The vector core consists of 16 single instruction, multiple data (SIMD) lanes and supports different matrix and vector sizes.
-The considered processing times for different matrix and vector operations are shown in Table I with $N=16$ and $K=16$ or $K=128$, together with the complexities order of each operation.
-The inverse of the Gramian is calculated using extended QRD, due to its high performance on this specific vector core.
+The vector core consists of 16 single instruction, multiple data lanes and supports different matrix and vector sizes.
+The processing times for different matrix and vector operations are shown in Table I with $N=16$ and $K=16$ or $K=128$, together with the complexities order of each operation.
+The inverse of the Gramian is calculated using extended QRD, due to its high performance on this specific vector core[9].
 The remaining digital logic blocks are implemented on the AMD ZCU216 FPGA to extract the latency numbers from real hardware.
@@
 This removes the need for long buffers in relation to the Ethernet.
-For all hardware units, a clock frequency of $150$MHz is assumed.
+For all hardware units, a clock frequency of 150150MHz is assumed.
 IV Latency Analysis
@@
 1 a break-down of the system latency is displayed.
-We define the latency as the time starting when the UE has new data to transmit and ending when the data for the first subcarrier has been detected by the BS, using either ZF or MRC algorithms.
+We define the latency as the time starting when the UE has new data to transmit and ending when the data for the first subcarrier has been detected by the base station (BS), using either ZF, MRC or MMSE algorithms.
 The latency can be expressed as
-$\tau=\tau_{\mathrm{wait}}+\tau_{\mathrm{LPU}}+\tau_{\mathrm{fronthaul}}+\tau_{\mathrm{CPU}}.$ (5)
+$\tau=\underbrace{\tau_{\mathrm{wait}}}_{\text{frame}}+\underbrace{\tau_{\mathrm{LPU}}+\tau_{\mathrm{fronthaul}}+\tau_{\mathrm{CPU}}}_{\text{PHY layer processing}}.$ (5)
 It consists of two main sources, the latency due to the frame structure $\tau_{\mathrm{wait}}$ and the latency due to the physical layer processing in the BS.
 Based on the architecture in Fig.
-1, the BS processing latency has three parts $\tau_{\mathrm{LPU}}$, $\tau_{\mathrm{fronthaul}}$, and $\tau_{\mathrm{CPU}}$, which are the latency in the BS cause by the LPU, front-haul and CPU respectively.
-A few assumptions are made regarding the system throughput to simplify latency calculations.
-It is assumed that there are enough hardware resources to match the data throughput of both in-panel processing and inter-panel data transfer.
-These assumptions gives the liberty to measure the individual latency of every source in the system without taking the throughput into account (e.g., the extra latency due to buffering can be neglected).
-There are also a few latency sources that we disregard in the analysis, as they are too small to have any impact.
-The latency due to the RF chain in the LPU is known to be very small[11] and, as it is done in parallel for all antennas, it can safely be ignored.
-Also, as an indoor scenario is considered, the signal propagation, both the one over air between the UE and BS, and the one between the panels, are not taken into account.
+2, the BS processing latency has three parts $\tau_{\mathrm{LPU}}$, $\tau_{\mathrm{fronthaul}}$, and $\tau_{\mathrm{CPU}}$, which are the latency caused by the LPU, fronthaul and CPU, respectively.
+The following assumptions are made to simplify latency calculations.
+Firstly, there are enough hardware resources to match the data throughput of both in-panel processing and inter-panel data transfer.
+This makes it possible to measure the individual latency of every source in the system without taking the throughput into account (e.g., the extra latency due to buffering can be neglected).
+Secondly, latency due to the RF chain in the LPU is known to be very small[10] and, since it is done in parallel for all antennas, it is ignored.
+The signal propagation, both the one over air between UE and BS, and the one between panels, scale linearly with distance traversed.
+Propagation over air takes about $3.3\mu s$ per km, and, with optical cables used, the total propagation between the panels for a daisy chain topology takes about $5.6\mu s$ per km of total distance traversed between all panels.
+As an indoor scenario is considered, the latency caused by these two steps is negligible and not taken into account.
+Even for an outdoor scenario, the distances need to be in the order of several km for it to have any significant impact.
 When computing $\tau_{\mathrm{wait}}$, we calculate the longest time from when the UE has new data to transmit until there is an available data (ULD) slot.
-This latency will depend on both the frame structure and the OFDM transmission time.
+This latency depends both on the frame structure and the OFDM transmission time.
 In this analysis, we consider a fixed frame structure as presented in Fig.
-2.
-It is composed of a total of 7 OFDM slots, consisting of UL and downlink (DL) transmission, as well as guardband.
-The OFDM transmission time, including the cyclic prefix, is $19\mu s$, and a subcarrier spacing of 60kHz is assumed.
-The highest $\tau_{\mathrm{wait}}$ happen when the UE gets new data to transmit as the second ULD symbol has just started transmitting, as it then has to wait for six OFDM symbols before the next available ULD slot.
+1.
+It is composed of a total of 7 OFDM slots, consisting of UL and downlink (DL) transmission, as well as a guard time for time-division duplexing switching.
+The OFDM transmission time, including cyclic prefix, is $19\mu s$, at a subcarrier spacing of 60kHz.
+The highest $\tau_{\mathrm{wait}}$ happens when the UE gets new data to transmit immediately after the second ULD symbol transmission has started, as it then has to wait for six OFDM symbols before the next available ULD slot.
 This makes $\tau_{\mathrm{wait}}=133\mu s$.
-Due to this long wait at the UE, any processing necessary at the UE side can safely be disregarded, as it will be able to finish within this time frame.
-Figure 3: The latency results, in both percentage and actual values, depending on the amount of UEs, for a system with 16 antennas per panel and 127 panels, for both ZF and MRC.
+Due to this long wait at the UE, any processing necessary at the UE side can safely be disregarded, as it will be able to finish within this time frame due to its simplicity.
 For LPU the latency, $\tau_{\mathrm{LPU}}$, is divided into two sources.
-The latency from the implemented digital front end and FFT block has a latency of $25\mu s$.
+The latency from the implemented digital front-end and FFT block has a latency of $25\mu s$.
 It is quite high due to the fact that the full OFDM symbol must be received before the computations can be completed.
@@
 Also, both the processing for the UL data symbol and the UL pilot need to be taken into account, as the calculations for the UL data cannot start until the channel estimate for the relevant subcarrier is available.
-In ZF, the latency for the local Gramian computation also needs to be taken into account, as it needs to be transmitted through the fronthaul architecture before the local MRC vector.
+In ZF and MMSE, the latency for the local Gramian computation also needs to be taken into account, as it also needs to be transmitted through the fronthaul architecture.
 The complexity of the relevant calculations are shown in Table I.
-The latency of the Ethernet-based fronthaul, $\tau_{\mathrm{fronthaul}}$, grows linearly with the amount of panels according to
+The latency of the Ethernet-based fronthaul, as implemented in the reference architecutre, $\tau_{\mathrm{fronthaul}}$, grows linearly with the amount of panels according to
 $\tau_{\mathrm{fronthaul}}=0.87\times(P-1),$ (6)
@@
 The complexity of these calculations is also included in Table I.
+Assuming the central unit knows the SNR the only difference between ZF and MMSE is an addition between the Gramian and a diagonal matrix, which is computed using $K$ additions.
+As this operation is both of much lower complexity, and consists of additions instead of multiplications, compared to the rest of the central processing, we can ignore the small extra latency this causes and assume that $\tau_{\mathrm{CPU}}$ for ZF and MMSE are the same.
+Figure 3: The latency, in both absolute and relative values, depending on the amount of UEs, for a system with 16 antennas per panel and 127 panels.
+Figure 4: Example scenarios used in the spectral efficiency simulations. $P=1$ with a $32\times 32$ panel (left) and $P=32$ with $8\times 4$ panels (right).
 IV-B Latency Results
-Figure 4: Example scenarios used in the spectral efficiency simulations. $P=1$ with a $32\times 32$ panel (left) and $P=32$ with $8\times 4$ panels (right).
 In Fig.
-3 the latency for a system with $N=16$ and $P=128$ is shown for both ZF and MRC.
-When the number of UEs is relatively small, the latency distribution of different components are similar, with almost half being due to the frame structure, roughly a third due to the Ethernet-based fronthaul and the rest caused by the processing.
-However, when the amount of UEs grows, the two approaches lead to completely different latency ratios.
-For MRC, increasing the amount of UEss does not affect the latency much, as the needed computations only grow linearly with $K$ and are quite simple to execute.
-For ZF though, the processing latency is completely dominant for a large amount of UEs as the local processing grows quadratically with $K$ and the central processing grows cubically with $K$.
-To decrease the latency of a distributed massive multiple-input multiple-output (DMIMO) system using ZF and which is designed to support a large amount of UEs, further optimizations clearly need to be done.
-A simple solution would be to further optimize the processing, though that would likely lead to higher resource and energy costs.
-Another interesting avenue that could be investigated is the potential of panel segmentation, where only a subset of panels would serve certain UEs.
-This would lead to smaller matrix sizes and therefore simplifying the processing, but it would cause new unique problems.
+3 the latency (absolute and relative) for a system with $N=16$ and $P=128$ is shown for ZF, MMSE and MRC.
+When the number of UEs is small, latency distributions are similar, with almost half of the latency due to the frame structure, roughly a third due to the Ethernet-based fronthaul, and the rest caused by the processing.
+However, when the amount of UEs grows, two completely different latency ratios appear.
+For MRC, increasing the amount of UEs does not affect the latency much, as the needed computations only grow linearly with $K$ and are simple to execute.
+For both ZF and MMSE though, central processing latency dominates at large numbers of UEs.
+While local processing grows quadratically with $K$, the central processing grows cubically with $K$.
+To decrease the latency of a distributed MIMO system, using ZF or MMSE, which is designed to support a large number of of UEs, further optimizations are clearly needed.
+A simple solution would be to further optimize the central processing latency, specifically the matrix inversion, though that would likely lead to higher resource and energy costs.
+Another interesting avenue is the potential of panel segmentation, where only a subset of panels serve certain UEs.
+This would lead to smaller matrix sizes and, therefore, simplifying the processing, but it would, in turn, most likely cause new problems.
 V Latency Versus Spectral Efficiency
 One of the main performance metrics in any wireless communication scenario is the data rate supported by the system.
-In the considered setting, we are specially interested in the data rates experienced by the different UEs when transmitting their UL data, and how they interplay with the arrangement of the LIS into panels, as well as with the resulting physical layer latency.
-To this end, we consider an industrial indoor scenario, where the channel matrix follows the Rician fading model [7]
+For the considered setting, we are specially interested in the data rates experienced by the different UEs in the UL, how these interplay with the arrangement of the LIS into panels, and the resulting physical layer latency.
+To this end, we consider an industrial indoor scenario, where the channel matrix
 $\boldsymbol{H}=\sqrt{\frac{K_{\mathrm{F}}}{1+K_{\mathrm{F}}}}\boldsymbol{H}_{\mathrm{LoS}}+\frac{1}{\sqrt{1+K_{\mathrm{F}}}}\boldsymbol{H}_{\mathrm{R}},$ (7)
-where $K_{\mathrm{F}}$ is the Rician factor, $\boldsymbol{H}_{\mathrm{LoS}}$ is the line of sight (LoS) channel between the UEs and the whole LIS, and $\boldsymbol{H}_{\mathrm{R}}$ follows the IID Rayleigh fading model where the average power for each entry equals the power of the respective LoS contribution in $\boldsymbol{H}_{\mathrm{LoS}}$.
+is assumed to follow a Rician fading model [7], with $K_{\mathrm{F}}$ being the Rician factor, $\boldsymbol{H}_{\mathrm{LoS}}$ the line of sight (LoS) channel between the UEs and the entire LIS, and $\boldsymbol{H}_{\mathrm{R}}$ an IID Rayleigh fading model where average power for each entry equals the power of the respective LoS contribution in $\boldsymbol{H}_{\mathrm{LoS}}$.
 The parameters of the scenario are summarized in Table II.
-Note that, since we consider LIS scenarios with a large number of panels deployed throughout compact indoor regions, we should take into account near-field effects when computing the LoS contributions [12].
+Note that, since we consider LIS scenarios with a large number of antenna elements deployed throughout compact indoor regions, we should take into account near-field effects when computing the LoS contributions [11].
 Specifically, we consider that only the perpendicular incident power is absorbed by each antenna element, and we construct the array steering vectors using the spherical propagation model.
@@
 Antenna spacing $\lambda/2$
-Rician factor $5$ dB
+Rician factor 55 dB
 Parameter Value
@@
 Room length $30$ m
-Room width $20$ m
+Room width 2020 m
 We begin by analyzing the effect of distributing a fixed number of antennas into a varying number of LIS-panels.
@@
 Note that, in all cases (except when all the antennas are co-located), each panel is facing another panel.
-The number of panels is incremented from $P=1$ to $P=256$ by powers of 2, while for each $P$ value the respective panels are either squared or have twice as many elements along their width than along their height.
-There are $K=128$ UEs transmitting equal power, and these are randomly located throughout a plane at a minimum distance of $10\lambda$ from the walls.
+The number of panels is incremented from $P=1$ to $P=256$ by powers of 2, while for each $P$ value the respective panels are either square or have twice as many elements along their width than along their height.
+There are $K=128$ UEs transmitting equal power, and these are randomly located on a plane at a minimum distance of $10\lambda$ from the walls.
 The LIS-panels extend from the UEs plane towards the ceiling of the room.
-An example scenario is given in Fig.
+Two example scenarios are shown in Fig.
 4.
-All the spectral efficiency results are averaged over $10^{2}$ realizations of the uniform UE placements, each including $10^{3}$ realizations of the Rayleigh fading channel component.
+All spectral efficiency results are averaged over $10^{2}$ realizations of the uniform UE placements, each including $10^{3}$ realizations of the Rayleigh fading channel component.
 Fig.
 5 plots the spectral efficiencies with respect to the number of panels for the described scenarios.
-We can see that distributing the antennas throughout the scenario into smaller LIS-panels favors user fairness, reducing potential outages for both ZF and MRC processing.
-We believe that this is a strong motivation for favoring more distributed antenna deployments since it increases the overall reliability.
-On the other hand, as seen in the previous section, in dense user deployments the latency values for the ZF case are dominated by the central processing required to perform matrix inversion, so the most distributed implementation has marginal latency increase over having all the antennas co-located in one panel.
+We can see that distributing the antennas throughout the scenario into more smaller LIS-panels favors user fairness, reducing potential outages for all processing schemes, while it also allows for a slight increase in average user rate.
+We believe that this is a strong motivation for favoring more distributed antenna deployments since it increases the overall system reliability.
+At the same time, as seen in the previous section, in dense user deployments the latency values for the ZF and MMSE cases are dominated by the central processing required to perform matrix inversion, so the most distributed implementation only has marginal latency increase over having all the antennas co-located in one panel.
 In the MRC case, the fronthaul latency associated to sharing the data sequentially between panels still has noticeable impact, leading to a latency-performance trade-off which should be further analyzed under different settings.
-Nevertheless, there is a major latency reduction with respect to ZF, at the cost of some degradation in average user spectral efficiency.
+Nevertheless, there is a major latency reduction with respect to ZF and MMSE, at the cost of some degradation in average user spectral efficiency.
 Figure 5: User spectral efficiencies versus number of panels in logarithmic scale for fixed total number of antennas.
 Figure 6: Average user spectral efficiency versus number of $4\times 4$ panels in logarithmic scale. For each point, the computed latency value is included.
-We now consider the scaling of the spectral efficiencies when deploying varying number $P$ of equally sized square panels with $N=16$ antennas.
+We now consider the scaling of the spectral efficiencies when deploying varying numbers, $P$, of equally sized square panels with $N=16$ antennas each.
 Fig.
@@
 Aside from the fixed panel size, the same scenario considerations apply as for the results in Fig.
-6.
-As one may expect, the average user spectral efficiency increases logarithmically with the number of panels for high enough $P$.
-This may be justified by the extra array gain associated to having more panels receiving the transmitted power.
-We have also included for comparison the latency values for each panel deployment.
-Interestingly, the spectral efficiency achieved by MRC is comparable to that achieved by ZF with half the LIS-panels, which still requires a latency close to $30$ times higher.
-Thus, we can conclude that, in the considered scenario, MRC is much more effective at exploiting the trade-off between latency and spectral efficiency.
+5.
+For large enough $P$, ZF and MMSE exhibit similar performance, hinting that such scenario is interference limited, while the average user spectral efficiency increases logarithmically with the number of panels, which is justified by the extra array gain associated to having more panels receiving the transmitted power.
+Fig.
+6 also includes the associated latency values for each panel deployment.
+Interestingly, the spectral efficiency achieved by MRC is comparable to that achieved by ZF and MMSE with a quarter of the LIS-panels, which still requires a latency close to $30$ times higher.
+Thus, we conclude that, in the considered scenario, MRC is much more effective at exploiting the trade-off between latency and spectral efficiency.
 VI Conclusion
 We have studied the latency in a panel-based LIS indoor scenario with different panel deployments.
-We presented algorithms for performing decentralized MRC and ZF processing without any loss with respect to the centralized counterpart.
-We also computed latency values based on real implementations for the different schemes, which show that, as the number of users grows, the ZF latency is quickly dominated by the central processing required to invert the channels.
-Finally, we compared the spectral efficiency achieved with different deployments, where we observed that having antennas more distributed increases the system reliability, and comes at essentially no latency cost when using ZF.
+We presented algorithms for performing decentralized MRC, ZF and MMSE processing without any loss compared to the centralized counterpart.
+We also compute latency values based on real implementations for the different schemes, which show that, as the number of users grows, the ZF and MMSE latency is quickly dominated by the central processing required to invert the channel Gramian.
+Finally, we compared the spectral efficiency achieved with different deployments, where we observed that having antennas more distributed increases the system reliability, and comes at essentially no extra latency cost when using ZF or MMSE.
 However, MRC seems to attain a better trade-off between latency and performance.